OpenStudy (anonymous):

can some one help me with this question. Four microprocessors are randomly selected from a lot of 100 microprocessors among which 10 are defective. Find the probability of obtaining no defective microprocessors

OpenStudy (anonymous):

90 choose 4 / 100 choose 4 = 0.651630549

OpenStudy (anonymous):

ok that was wrong

OpenStudy (anonymous):

^ choosing 4, not 5 mang

OpenStudy (anonymous):

no doubt the same thing marine wrote, but my original answer was wrong

OpenStudy (anonymous):

\[\frac{90}{100}\times \frac{89}{99}\times \frac{88}{98}\times \frac{87}{97}\]

OpenStudy (anonymous):

yeah wrong twice!